OpenStudy (teamus1021):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = 3x + 10 y = -3x - 2

OpenStudy (3mar):

May I help?

OpenStudy (teamus1021):

y = 1/3 x + 14/3 y=-1/3 x +10/3

OpenStudy (teamus1021):

@jigglypuff314

OpenStudy (teamus1021):

sure @3mar

OpenStudy (3mar):

Thanks. For parallel lines: "All parallel lines have the same slope!" So the idea is to find the slope of this line (y = -3x + 6) and it is the same as the slope of your line you want to form its equation. Now use this slope and the point you are given to form it! http://pasteboard.co/FwTzngLl.png I hope I helped you.

OpenStudy (teamus1021):

i dont understand

OpenStudy (3mar):

Ok. Step by step. What is the slop of the line (y = -3x + 6)?

OpenStudy (teamus1021):

2

OpenStudy (mathmale):

All right. What next? You have the slope, m, and one point that lies on the line.

OpenStudy (mathmale):

Which would be the most straightforward equation of a line to use here, with that info?

OpenStudy (teamus1021):

i dont know anything about algebra this why im up here for help and understanding

OpenStudy (3mar):

To find the slope of a line, put it in the form of the slope-intercept form.